OpenStudy (anonymous):

What is the period of vibration for a wave with a frequency of 150 hertz? a. 0.006 seconds b. 0.06 seconds c. 15 seconds d. 150 seconds Can you help me solve it

OpenStudy (anonymous):

frequency =1/period

OpenStudy (anonymous):

so period = 1/frequency

OpenStudy (anonymous):

\[\frac{ 1 }{ 150 } = 6.667 \times 10^{-3}\]

OpenStudy (anonymous):

This is our formula: Period (T) = 1/Frequency (f)

OpenStudy (ladiesman217):

We use this equation: Period (T) = 1/Frequency (f) Basically, we substitute the values in to solve for period (T): T = 1/f T = 1/(150) what does t equal after dividing 1 and 150

OpenStudy (anonymous):

its a. :D

OpenStudy (anonymous):

am i rite

OpenStudy (ladiesman217):

yes you are good job

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Thank you guys <3 you guys are the best

OpenStudy (ladiesman217):

your welcome

OpenStudy (anonymous):

yw